In Hawaii shark attacks are very rare. Furthermore, about 55% of all shark attacks are serious, but not fatal.† You are scheduled to give a talk at a local scuba club in Honolulu and want to include a brief statistical profile regarding the most recent nine shark attacks, all of which were nonfatal. Let r be the number of nonfatal attacks out of a random sample of nine shark attacks. (a) Make a histogram for the probability distribution of r. WebAssign Plot WebAssign Plot WebAssign Plot WebAssign Plot (b) What is the probability all nine shark attacks are nonfatal? (Round your answer to three decimal places.) (c) What is the probability that seven or more of the nine shark attacks are nonfatal? (Round your answer to three decimal places.) (d) What is the expected number of nonfatal shark attacks out of the nine? What is the standard deviation of the r-probability distribution? (Round your answer to three decimal places.)
Answer #1

a)

b)

P(X=9)= (nCx)px(1−p)(n-x)    = 0.005

c)

P(X>=7)=1-P(X<=6)= 1-∑x=0x-1   (nCx)px(q)(n-x) = 0.150

d)

expected value E(x)=μ=np=4.95
standard deviation σ=√(np(1-p))=1.492